OpenStudy (anonymous):

If cos x= -3/4, find cos 2x

OpenStudy (anonymous):

\[\cos(2x)=2\cos^2(x)-1\] so \[\cos(2x)=2(-\frac{3}{4})^2-1\]

OpenStudy (anonymous):

10